Find all solutions of the given differential equations and then find the particular solutions for which a point (x,y) is given:


dy/dx = √x + 3 ; (x,y) =(-1,3)


1
Expert's answer
2021-09-02T14:26:22-0400

First we solve the integral by separating the terms:


"\frac{dy}{dx}=\sqrt{x}+3 
\\ \implies \intop dy=\int(\sqrt{x}+3)dx
\\y=\int\ x^{1/2}dx+3\int dx=\frac{x^{3/2}}{3/2}+3x+C
\\y=\frac{2x^{3/2}}{3}+3x+C"


Then, we substitute the coordinates (x,y) to find C for the particular solution:


"3=\frac{2(-1)^{3/2}}{3}+3(-1)+C 
\\ \implies 3=-\frac{2i}{3}-3+C
\\ \to C=6+\frac{2i}{3}"


In conclusion:


"\text{General solution: }
\\ y=\frac{2x^{3/2}}{3}+3x+C
\\ \text{Particular solution: }
\\ y=\frac{2x^{3/2}}{3}+3x+6+\frac{2i}{3}"